OpenStudy (anonymous):

Use the functions given by f(x) = (-x/8)+1 and g(x) = x^3 to find (f o g)^-1 (-5).

OpenStudy (xapproachesinfinity):

well what is your fog? Did you do that first

OpenStudy (anonymous):

I am stuck between (f(g(x)))^-1 or (f^-1(g^-1(x)))

OpenStudy (xapproachesinfinity):

Of course they are asking the first one But firstly, did you find f(g(x))

OpenStudy (anonymous):

Isn't it: (f(g(x)))^-1 = (f(x^3))^-1 = (((-x^3)/8)+1)^1

OpenStudy (xapproachesinfinity):

why are you doing it that way! you are mixing things that should not be mixed First find fog! forgot about inverse thing for a moment get the function.

OpenStudy (anonymous):

(f o g) = [(-x^3)/8] +1

OpenStudy (xapproachesinfinity):

Good! Now since that function is 1 to 1 then it has an inverse let's call that function h(x)=(fog)(x) from inverse property we now that if \(h^-1(-5)=a,~ then~ h(a)=-5\) then we have \(\Large\frac{-a^3}{8}+1=-5\) solve for a

OpenStudy (xapproachesinfinity):

know that*

OpenStudy (xapproachesinfinity):

that's \(h^{-1}(-5)=a\) in 4th line (correction)

OpenStudy (anonymous):

So, (-x^3)/8 = -6 -x^3 = -48 x^3 = 48 x = cube root of 48

OpenStudy (xapproachesinfinity):

yes! simplify that more

OpenStudy (anonymous):

I mean x is a, and a = 2 * cube root of 6

OpenStudy (xapproachesinfinity):

that's okay a or x doesn't matter, it's just a matter of choosing a variable so now we go back to our line above we said \(h^{-1}(-5)=a\\ so~~ h^{-1}(-5)=2\sqrt[3]{6}\)

OpenStudy (xapproachesinfinity):

and our h^-1(x) is just (fog)^-1(x) as we set it

OpenStudy (anonymous):

Ok, this makes so much more sense now! Thank you for all of your help.

OpenStudy (xapproachesinfinity):

welcome^_^